OpenStudy (anonymous):

y varies jointly with p and q. When p = 12 and q = 4, y = 144. Find y when p = 10 and q = 6.

OpenStudy (anonymous):

y = kpq 144 = k(12)(4) k = 3 y = 3(10)(6) y = 180 :)

OpenStudy (anonymous):

Thank you, again :) lol

OpenStudy (anonymous):

your welcome :))

OpenStudy (radar):

Calculate the constant k y=kpq 144=(12)(4) k k=3 y=(3)(10)(6)=180

OpenStudy (radar):

very good kumarsajan, the step by step really helps the student.

OpenStudy (anonymous):

thank you radar. :))

OpenStudy (jamesj):

Agreed.